OpenStudy (anonymous):

Factor: 2a^2 + 30a + 100

Parth (parthkohli):

A good start is to factor out the 2. \( \color{Black}{\Rightarrow 2(a^2 + 15a + 50)}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2(a^2 + 10a + 5a + 50)}\) Can you manage now? Yes?

OpenStudy (anonymous):

i got that part, what what's next?

OpenStudy (anonymous):

i think its 50

OpenStudy (jiteshmeghwal9):

First breackout 30a such that if we multiply them we get 100a^2 & if we add them we get 30a

OpenStudy (anonymous):

2a^2 + 30a + 100 factor completely 2(a^2 + 15a + 50) 2(a + 10)(a + 5) works as 10a + 5a = 15a & 10*5 = 50

OpenStudy (anonymous):

Thank youuu!

OpenStudy (jiteshmeghwal9):

so,\[2a^2+20a+10a+100\]\[2a(a+10)+10(a+10)\]\[(2a+10)(a+10)\]

Parth (parthkohli):

Hmm.

OpenStudy (anonymous):

wait, which one is right?

Parth (parthkohli):

@elitepenguin Please don't \(\textbf{answer}\) the questions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2a^2+30a+100 2(a^2+15a+50) 2(a^2+10a+5a+50) 2(a{a+10}+5{a+10}) 2({a+10}{a+5})

OpenStudy (anonymous):

basically you have to break the middle term (i.e 15a)into two values in such a way that when u multiply them u get 50 and when you add or minus them you get plus 15 for this question we use 10 and 5 10*5=50 10+5=+15